OpenStudy (anonymous):

Help!!! Find the equation of y, given the derivative and the indicated point on the curve? dy/dx = 2(x-1), point (3, 2)

OpenStudy (anonymous):

first expand the equation into 2x-2. Now integrate it with respect to x, you should end up with: x^2-2x+C where C is a constant

OpenStudy (amistre64):

we need to make sure the copy and paste didnt eat an exponent marker

OpenStudy (binary3i):

dy/dx=2(x-1) dy=2(x-1)dx integrate this. integration will give you a function(F(x)) of x and a constant y= F(x) + C so put the point in the place of X and Y and get your C do it

OpenStudy (anonymous):

Now substitute the point (3,2) into it 2=(3)^2-2(3)+C and solve for C Then you should arrive at your final equation

OpenStudy (anonymous):

Where should I substitute the point (3, 2)?

OpenStudy (anonymous):

Oh! I think I got it! Thank you so much!

OpenStudy (anonymous):

So, should it be y = x^2 - x -1?